a bodu of mass m1 , moving with velocity u1 collides head  on with a body of mass m2 moving with a velocity u2 the two bodies stick together after the collision the loss of kinetic energy after the collision is ?

Grade:

4 Answers

Nimit Yadav
11 years ago

Since the objects stick together they will be a part of plastic collision...

so. conservation of kinetic energy cannot be applied...

apply conservation of linear momentum and then find individual velocities

then loss of kinetic energy will be change in kinetic energy......

So,it will become equal to amount of work done

chhers you will get the answer

Prajwal kr
11 years ago

Here are a few ideas for you to build upon:

 

Conserve momentum:

m1u1 + m2u2 = (m1 + m2 )v............................. v is final velocity.

 

Now you know the final velocity. Calculate the final knetic energy and initial kinetic energies and find their change.

Akash Kumar Dutta
11 years ago

Dear Saurabh,

initial KE of the system= 1/2 [ m1.u1^2  +  m2.u2^2 ]
applying conservation of momentum we get new velocity of the new body=  (m1.u1-m2.u2) / (m1+m2)
new KE of the Body=      [  (m1.u.1 - m2.u2) / 2.(m1+m2)   ]^2
Subtracting and simplyfing we get the loss in KE after Collision=  [ m1.m2. (u1 + u2)^2 ] / [ 2.(m1 + m2) ]
Hope I Helped You.

Regards

                                              

Rohan Das
11 years ago

using conservation of momentum calculate final velocity.using K.E.=mv*v/2,we will get K.E.